OpenStudy (anonymous):

Order from greatest to least a) 25^100 b) 2^300 c) 3^400 d) 4^200 e) 2^600

OpenStudy (anonymous):

2^600, 3^400 25^100, 4^200, 2^300

OpenStudy (callisto):

Explanation?

OpenStudy (anonymous):

Yeah I know already. 25^100 2^300 = (23)100 = 8^100 3^400 = (34)100 = 81^100 4^200 = (42)100 = 16^100 2^600 = (26)100 = 64^100

OpenStudy (anonymous):

kk thanks :p

OpenStudy (tester97):

Want to know a easy way to solve this? take away the extra zero's a) 25^1 b) 2^3 c) 3^4 d) 4^2 e) 2^6 Now you simplify that. a)25 b)8 c)81 d)16 e)64 Now put that in order that should be easy for you yes?

OpenStudy (anonymous):

Oooooh. I hate zeros >_> Thank you :)

OpenStudy (tester97):

I explained this correct, yes? @Callisto